### v1

@@ 6 Conclusions and Future Work
In this paper, we presented an efficient algorithm for the APSP problem for a dynamic set of strings $\mathcal{S}=\{S_{1},\ldots,S_{k}\}$ of total length $n$. Our algorithm occupies $O(n)$ space and takes $O(|S_{i}|\log\sigma+i)$ time for each string $S_{i}$ inserted, for $1\leq i\leq k$. The running time is near-optimal except for the $\log\sigma$ factor, that is indeed required if one maintains $\mathcal{S}$ in a tree or DAG based data structure for a general ordered alphabet of size $\sigma$.
An intriguing question is whether our method can be modified to handle deletions of strings from $\mathcal{S}$. Diptarama et al. [11] showed how to update the failure links of the AC-automaton in $O(\sigma|S|+u_{f})$ time when a string $S$ is deleted from $\mathcal{S}$, where $u_{f}$ is the number of failure links that need to be updated. It is interesting to see whether our DAWG-based method can be modified for the fully dynamic set of strings allowing for both insertions and deletions.
It is also interesting to see whether our dynamic algorithm can be extended to efficient construction/update of the HOG [6, 12, 19] for a dynamic set of strings.
{credits}
6.0.1 Acknowledgements
This work was supported by JSPS KAKENHI Grant Numbers and JP20H05964, JP23K24808, JP23K18466 (SI).

### v2

@@ 7 Conclusions and Future Work
In this paper, we presented an efficient algorithm for the APSP problem for a dynamic set of strings $\mathcal{S}=\{S_{1},\ldots,S_{k}\}$ of total length $n$. Our algorithm occupies $O(n)$ space and takes $O(|S_{i}|\log\sigma+k)$ time for each string $S_{i}$ inserted or deleted. The running time is near-optimal except for the $\log\sigma$ factor, that is indeed required if one maintains $\mathcal{S}$ in a tree or DAG based data structure for a general ordered alphabet of size $\sigma$.
It is interesting to see whether our (fully) dynamic algorithm can be extended to efficient construction/update of the HOG [6, 12, 22] for a (fully) dynamic set of strings.
Acknowledgements
This work was supported by JSPS KAKENHI Grant Number JP23K24808 (SI).